OpenStudy (anonymous):

What is the arc length of x^(3/2) on the interval [2,3]... having trouble any help ?

OpenStudy (anonymous):

\[\int\limits_a^b \sqrt{1+f'(x)^2} \, dx\] f'(x)=\[\frac{3 \sqrt{x}}{2}\] \[f'(x)^2=\frac{9 x}{4}\] \[\int\limits _2^3\sqrt{1+\frac{9 x}{4}}\text{dx}\]

OpenStudy (anonymous):

\[\int\limits _2^3\sqrt{\frac{(4+9 x)}{4}}\text{dx}\]

OpenStudy (anonymous):

\[\frac{1}{2}\int\limits _2^3\sqrt{(4+9 x)}\text{dx}\]

OpenStudy (anonymous):

shouldnt that be (1/4)

OpenStudy (anonymous):

nevermind

OpenStudy (anonymous):

still cant

OpenStudy (anonymous):

Just use U-Sub